OpenStudy (anonymous):

what are the next two terms? 31,24,17,...

OpenStudy (anonymous):

here are the answer choices 1/2, 9/2 10,3, 13/2, 7 -14,-22 17/3, 7 -12, -23 22, 26 15/3, 20/3

OpenStudy (anonymous):

10,3

OpenStudy (anonymous):

how do you know?

OpenStudy (anonymous):

you subtract 7 each time 24-7=17-7=10-7=3

OpenStudy (anonymous):

oh ohk thanks :)

OpenStudy (anonymous):

np